OpenStudy (anonymous):

please help... Find an equation in standard form for the hyperbola with vertices at (0, ±2) and foci at (0, ±11).

OpenStudy (anonymous):

@openstudier101x

OpenStudy (anonymous):

somebody... anybody please....

OpenStudy (anonymous):

@amistre64 pleeeease...

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Loser66

OpenStudy (amistre64):

its a hyperbola, whats so tough about it?

OpenStudy (anonymous):

i just don't get precalc... .-.

OpenStudy (amistre64):

give me something to work with, what would you start with?

OpenStudy (anonymous):

i don't even know how to begin this problem....

OpenStudy (anonymous):

don't hate me for being annoying...

OpenStudy (amistre64):

well we know its a hyperbola, so lets start with the equation for one ....

OpenStudy (anonymous):

(x-h)^2/a^2 - (y-k)^2/b^2=1

OpenStudy (amistre64):

good, at worst we might have to swap x and y parts ... now tell me what you think h and k represent on our construction

OpenStudy (anonymous):

h is +/- 2 and k is +/-11??

OpenStudy (amistre64):

not quite, h and k are generally used for center point of a function, it where we grab hold of it to move it about the grid. what is the center of our hyperbola? what point is the center between the verts?

OpenStudy (anonymous):

(0,0)?

OpenStudy (anonymous):

i'm telling you i don't know...

OpenStudy (amistre64):

yep, seems fair, so h and k are 0 so we can ignore them

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

x^2/a^2 - y^2/b^2 = 1 now plug in a vertex, its an x and a y value so lets use it to start adjusting this thing.

OpenStudy (anonymous):

x^2/+/-2 -y^2/ +/-11+1

OpenStudy (anonymous):

opps =1

OpenStudy (amistre64):

just one point (0,2) is fine 0^2/a^2 - 2^2 /b^2 = 1 this simplifies to: - 2^2/b^2 = 1 what do you spose b will be?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

meh, b^2 = 4, so b=2 now we have something to adjust - 2^2/2^2 = 1 gives us: -1=1 does this make any sense? or should y be positive instead?

OpenStudy (anonymous):

idk man.... is this the final answer?? y^2/121 - x^2/4 =1

OpenStudy (amistre64):

no, its not

OpenStudy (anonymous):

okay is it y^2/4 - x^2/121 =1??

OpenStudy (amistre64):

im not playing a guessing game, im taking my time trying to help you learn it .. if you dont want to take the time ill just go do something else.

OpenStudy (anonymous):

okay :/

OpenStudy (amistre64):

should i stay or go?

OpenStudy (anonymous):

stay

OpenStudy (amistre64):

ok, then we still have a question - 4/4 = 1 is telling us that -1=1 which is not right. would you agree that this means that the y parts need to be positive instead?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then lets adjust instead of: x^2 /a^2 - y^2 /4 = 1 we need to swap the x and y parts y^2/4 - x^2/a^2 = 1 now we have one part left to determine, and we can do that with the pythag thrm the distance a focus is from the center is a hypotenuse, a and b are legs. whats the distance to our focus?

OpenStudy (anonymous):

ummm.... 11^2 so 121..??

OpenStudy (amistre64):

well 11, but yeah a^2 + b^2 = 11^2 a^2 = 121 - 4

OpenStudy (anonymous):

so the final answer is Y^2/4-x^2/117=1?

OpenStudy (amistre64):

thatll do it

OpenStudy (anonymous):

thank you! :)

OpenStudy (amistre64):

youre welcome